# leftmargin=0pt, rightmargin=0pt, backgroundcolor=bggray, middlelinecolor=black, roundcorner=3 *[subfigure]position=bottom

### v1

@@ 6.2.2. Effects on Gender Disaggregated by Trustworthiness
At the Story Level
While we have seen that participants wrote significantly fewer masc-coded detective and significantly more fem-coded detective stories with fem-coded suggestions, we find in the disaggregated results (Figure 7) that this is only true when the partner detective is untrustworthy (untrustworthy: $t(113)=-4.367$, $p_{\text{FDR}}\approx 0.0003$; trustworthy: $t(119)=-1.561$, $p_{\text{FDR}}\approx 0.2634$). This happens because without suggestions, participants wrote significantly more masc-coded partners when they were untrustworthy ($t(71)=-4.117$, $p_{\text{FDR}}\approx 0.0008$) and significantly more fem-coded partners when they were trustworthy ($t(71)=3.051$, $p_{\text{FDR}}\approx 0.0161$). On the other hand, masc-coded suggestions led to significantly more untrustworthy masc-coded partners ($t(169)=-2.881$, $p_{\text{FDR}}\approx 0.0206$) and marginally more trustworthy fem-coded partners ($t(169)=2.307$, $p_{\text{FDR}}\approx 0.0735$), leading to no significant differences between stories written without suggestions or with masc-coded suggestions for any gender or trustworthiness.
These results show that control stories written without suggestions are more similar to those written with masc-coded detective suggestions, regardless of trustworthiness (H1). Further, we see that participants wrote or accepted untrustworthy masc-coded characters than trustworthy masc-coded characters. While fem-coded suggestions were not affected by trustworthiness, we also see that without suggestions, participants are more comfortable writing trustworthy fem-coded characters than trustworthy masc-coded characters. This is consistent with Cao et al. (2022b)’s finding that annotators tend to view women as more trustworthy than untrustworthy and men more untrustworthy than trustworthy.
@@
When we disaggregate by trustworthiness (Table 4), we surprisingly see significantly more newly written or edited words when suggesting a trustworthy fem-coded partner than untrustworthy ($t(3443)=-3.168$, $p_{\text{FDR}}\approx 0.0085$), and we only see a significant difference between genders in the trustworthy case ($t(3907)=5.235$, $p_{\text{FDR}}<0.0001$), with more new typing in the trustworthy fem-coded partner case. When we constrain to words that determine gender, the only remaining significant trend is the increased participant contribution in the trustworthy fem-coded partner case ($t(161)=-3.298$, $p_{\text{FDR}}\approx 0.0069$). This appears to go against our hypothesis that participants will rely more on pro-stereotypical suggestions than anti-stereotypical. However, when the participants decide to override the suggestions (by writing a new word or editing a suggestion), this analysis does not take into account what gender is being expressed in the override. While we observed double the overrides in the trustworthy fem-coded partner case than untrustworthy, we also observe that more of the overrides in the trustworthy case ultimately still produce a fem-coded partner ($27.4\%$) than in the untrustworthy case ($12.9\%$).
When we consider the rates of gender-specifying words being accepted or rejected, we see no significant difference between untrustworthy and trustworthy fem-coded suggestions ($t(515)=1.364$, $p_{\text{FDR}}\approx 0.3400$). Instead, we see significantly more acceptance of masc-coded suggestions than fem-coded regardless of trustworthiness (trustworthy: $t(507)=-4.224$, $p_{\text{FDR}}\approx 0.0003$; untrustworthy: $t(325)=-7.066$, $p_{\text{FDR}}<0.0001$) and significantly more acceptance of untrustworthy masc-coded suggestions than trustworthy ($t(317)=4.384$, $p_{\text{FDR}}\approx 0.0002$ ).
Overall, we see that the rejection rate results concur with Cao et al. (2022b)’s findings about gender-trustworthiness stereotypes and support H2b. While our results do not support H2 and H2a (about overall reliance and the source of gender-defining words), we note that these could be caused by a difference in the rate of overrides that change how/when the partner’s gender is expressed but not what that gender is.
@@ 6.3. Scenario: Wedding
In this scenario, we consider a wedding between two partners. We vary whether the partner who is mentioned in the story prefix is named “Joe” (a traditionally masculine name) or “Sherry” (a traditionally feminine name) and prompt the LLM to suggest that the second partner in the couple to be fem-coded or masc-coded (see examples in Table 5). We treat the pro-stereotypical conditions to be those where the genders of the partners are suggested to be different and the ani-stereotypical to be those where they are suggested to be the same.
@@
Gender breakdown of second partners with varied suggestions. When the first partner is named “Joe”, we see more fem-coded second partners (i.e. a straight relationship). We see more masc-coded partners with queer suggestions, but even with these suggestions, we still see more straight relationships than queer. When the first partner is named “Sherry”, we see the opposite effect. Here, we do see more acceptance of queer relationship suggestions than for “Joe”.
@@
We first focus on the rates of sexualities present in the overall stories (Figure 8). We start by considering what kinds of suggestions yield similar or different stories to those written without suggestions. When the first partner is “Sherry” and the model suggests a queer relationship (i.e., that Sherry’s partner is fem-coded), we see significantly fewer stories where the other partner is masc-coded ($t(132)=2.973$, $p_{\text{FDR}}\approx 0.0172$) and significantly more stories where the other partner is fem-coded ($t(132)=-2.695$, $p_{\text{FDR}}\approx 0.0328$). There are no significant changes to the distribution of Sherry’s partners’ gender when straight suggestions are provided (fem-coded partner: $t(124)=0.657$, $p_{\text{FDR}}\approx 0.7019$; masc-coded partner $t(124)=-0.422$, $p_{\text{FDR}}\approx 0.8234$). The direction of these trends are mirrored for “Joe” stories but with insignificant changes. That is, “Joe” is written with masc-coded partners more often ($t(115)=-1.688$, $p_{\text{FDR}}\approx 0.2218$) and fem-coded partners less often when given queer suggestions ($t(115)=2.143$, $p_{\text{FDR}}\approx 0.1071$), and there are no apparent changes when given straight suggestions to the rate of fem-coded ($t(104)=-0.556$, $p_{\text{FDR}}\approx 0.7514$) or masc-coded partners ($t(104)=-0.154$, $p_{\text{FDR}}\approx 0.9479$). These findings suggest that indeed participants’ default behaviour without suggestions is more similar to the “straight” conditions than the “queer” ones, though participants seem to be more resistant to accepting “queer” suggestions in the masc-coded case.
@@
We also consider within a suggestion type when the difference between choosing fem-coded and masc-coded partners is significant. When the first partner is named “Joe”, we see significantly more fem-coded partners than masc-coded regardless of the presence or type of suggestions (No suggestions: $t(56)=-8.825$, $p_{\text{FDR}}<0.0001$; queer suggestions: $t(174)=-6.035$, $p_{\text{FDR}}<0.0001$; straight suggestions: $t(152)=-16.063$, $p_{\text{FDR}}<0.0001$). For “Sherry”, we see significantly more masc-coded partners than fem-coded when participants see no suggestions ($t(86)=7.704$, $p_{\text{FDR}}<0.0001$) or straight suggestions ($t(162)=12.859$, $p_{\text{FDR}}<0.0001$). We see a trend in the same direction with queer suggestions, but here it is not significant ($t(178)=2.12$, $p_{\text{FDR}}\approx 0.1097$). These findings imply that without suggestions, participants default to hetero-normative stories and continue to write them when prompted (H1). When queer stories are suggested, depending on the gender of the first partner mentioned, participants either continue to prefer hetero-normative stories (in the case of “Joe”) or start to accept more queer stories (in the case of “Sherry”).
@@
Based on these results, even when writing with a “perfectly debiased” predictive text system that has no preference for gender pairings, we would expect to continue to see far more straight stories than queer ones.
@@
[…]
Gender and benevolence breakdowns for president stories. When considering gender alone, we see that masc-coded suggestions yield a similar breakdown as no suggestions (with much more masc-coded presidents than fem-coded). With fem-coded suggestions, we see more fem-coded presidents overall, but there are still more masc-coded presidents than fem-coded. When we also consider benevolence, we see that these suggestions were less effective. Where participants specified if the president is benevolent/threatening, they generally marked the president as benevolent, regardless of the presence or type of suggestions.
We continue by assessing the word-level acceptance and overriding of model suggestions (Table 4). When we constrain only to the words that affect the second partner’s gender and the pair’s inferred sexuality (H2a), we see that participants type their own gender-defining words more when the first partner is masc-coded regardless of whether the combination of genders match or do not (fem-coded queer vs masc-coded queer: $t(206)=-2.566$, $p_{\text{FDR}}\approx 0.0424$; fem-coded straight vs masc-coded straight: $t(195)=-3.554$, $p_{\text{FDR}}\approx 0.0031$). When we widen to all words written in the stories (H2), we continue to see a higher proportion of words coming from model suggestions in the conditions where the first partner is fem-coded (fem-coded queer vs masc-coded queer: $t(4230)=-4.019$, $p_{\text{FDR}}\approx 0.0005$; fem-coded straight vs masc-coded straight: $t(3699)=-4.373$, $p_{\text{FDR}}\approx 0.0001$), but now we also see we see marginally more acceptance of model suggestions in the condition where the fem-coded partner is suggested to be marrying a masc-coded character ($t(4010)=2.381$, $p_{\text{FDR}}\approx 0.0610$). Considering the rate of accepting vs rejecting gender-defining model suggestions (H2b), we further see that queer suggestions are more rejected for masc-coded characters than fem-coded ($t(546)=6.786$, $p_{\text{FDR}}<0.0001$) and that for masc-coded characters, queer suggestions are rejected significantly more than straight suggestions ($t(607)=-2.901$, $p_{\text{FDR}}\approx 0.0187$).
6.4. Scenario: President
In this scenario, the story prefix describes a new president laying out their policy vision (“In the first 100 days, the new president was determined to focus”). The model is prompted (unseen by the participant) to suggest that the president is of a particular gender (man or woman) and communion (“benevolent” or “threatening”). We provide example co-written stories with each character attribute in Table 5. The in-context examples given to the model focused on the benevolence axis as it applies to foreign policy, but the final written stories discuss a wider array of policy areas. According to Cao et al. (2022b), American annotators view men as comparatively more threatening than women.
6.4.1. Effects on Gender Alone
First, we consider how suggestions affect how participants specify a president character’s gender.
At the Story Level
@@ 6.4.1. Effects on Gender Alone
We first look into how suggestions affect the president character’s gender in the overall stories (9(a)). We find that there is no significant difference in the rate of making the president masc-coded without suggestions vs with masc-coded suggestions ($t(248)=-0.137$, $p_{\text{FDR}}\approx 0.9488$) and similarly for making the president fem-coded with masc-coded suggestions ($t(248)=-1.106$, $p_{\text{FDR}}\approx 0.4750$). However, when the model suggests a fem-coded president, participants write significantly fewer masc-coded presidents ($t(234)=-3.576$, $p_{\text{FDR}}\approx 0.0029$) and marginally more fem-coded presidents ($t(234)=2.429$, $p_{\text{FDR}}\approx 0.0565$) than those who did not receive suggestions. Further, when we compare the rates of making the president masc-coded vs fem-coded within conditions, we see that there are significantly fewer fem-coded presidents than masc-coded presidents in every condition, even with fem-coded suggestions (No suggestions: $t(146)=-8.947$, $p_{\text{FDR}}<0.0001$; masc-coded suggestions: $t(350)=-14.755$, $p_{\text{FDR}}<0.0001$; fem-coded suggestions: $t(322)=-4.343$, $p_{\text{FDR}}\approx 0.0002$).
@@
These results support H1, namely that the distribution of the gender of president characters when people write by default without suggestions is more similar to the distribution when people write with masc-coded suggestions than fem-coded suggestions. Based on these findings, we would expect stories written with a “debiased” predictive text model to still yield significantly more stories with masc-coded presidents than fem-coded presidents.
@@
Here we focus on word-level reliance on suggestions in gendered conditions (Table 3). We find participants’ overall reliance on model suggestions is not affected by the gender condition (H2; $t(6734)=-0.979$, $p_{\text{FDR}}\approx 0.5407$). However, when we only consider the story words that specify the president character’s gender, we see a lower rate of overrides or edits in the masc-coded president conditions (H2a; $t(203)=5.269$, $p_{\text{FDR}}<0.0001$). Similarly, when we consider only model suggestions that would specify the president character’s gender, we see a significantly higher rejection rate for words that would make the president fem-coded (H2b; $t(836)=6.362$, $p_{\text{FDR}}<0.0001$). These results support H2a-b as participants accept more model suggestions of masc-coded presidents than fem-coded presidents.

### v2

@@ 6.2.2. Effects on Gender Disaggregated by Trustworthiness
At the Story Level While we have seen that participants wrote significantly fewer masc-coded detective and significantly more fem-coded detective stories with fem-coded suggestions, we find in the disaggregated results (Figure 7) that this is only true when the partner detective is untrustworthy (untrustworthy: t(113)=4.367t(113)=-4.367, $p_{\text{FDR}}\approx 0.0003$, d=0.85d=-0.85; trustworthy: t(119)=1.561t(119)=-1.561, $p_{\text{FDR}}\approx 0.2634$, d=0.322d=-0.322). This happens because without suggestions, participants wrote significantly more masc-coded partners when they were untrustworthy (t(71)=4.117t(71)=-4.117, $p_{\text{FDR}}\approx 0.0008$, d=0.971d=-0.971) and significantly more fem-coded partners when they were trustworthy ($t(71)=3.051$, $p_{\text{FDR}}\approx 0.0161$, $d=0.72$). On the other hand, masc-coded suggestions led to significantly more untrustworthy masc-coded partners (t(169)=2.881t(169)=-2.881, $p_{\text{FDR}}\approx 0.0206$, d=0.445d=-0.445) and marginally more trustworthy fem-coded partners ($t(169)=2.307$, $p_{\text{FDR}}\approx 0.0735$, $d=0.356$), leading to no significant differences between stories written without suggestions or with masc-coded suggestions for any gender or trustworthiness.
These results show that control stories written without suggestions are more similar to those written with masc-coded detective suggestions, regardless of trustworthiness (H1). Further, we see that participants wrote or accepted untrustworthy masc-coded characters more than trustworthy masc-coded characters. While fem-coded suggestions were not affected by trustworthiness, we also see that without suggestions, participants are more comfortable writing trustworthy fem-coded characters than trustworthy masc-coded characters. This is consistent with human-held stereotypes as measured in our post-survey and in cao-etal-2022-theory—namely that people tend to view women as more trustworthy than untrustworthy and men more untrustworthy than trustworthy.
@@
When we disaggregate by trustworthiness (Table 4), we surprisingly see significantly more newly written or edited words when suggesting a trustworthy fem-coded partner than untrustworthy (t(3443)=3.168t(3443)=-3.168, $p_{\text{FDR}}\approx 0.0085$, d=0.108d=-0.108), and we only see a significant difference between genders in the trustworthy case ($t(3907)=5.235$, $p_{\text{FDR}}<0.0001$, $d=0.168$), with more new typing in the trustworthy fem-coded partner case. When we constrain to words that determine gender, the only remaining significant trend is the increased participant contribution in the trustworthy fem-coded partner case (t(161)=3.298t(161)=-3.298, $p_{\text{FDR}}\approx 0.0069$, d=0.52d=-0.52). This appears to go against our hypothesis that participants will rely more on pro-stereotypical suggestions than anti-stereotypical. However, when the participants decide to override the suggestions (by writing a new word or editing a suggestion), this analysis does not take into account what gender is being expressed in the override. While we observed double the overrides in the trustworthy fem-coded partner case as untrustworthy, we also observe that more of the overrides in the trustworthy case ultimately still produce a fem-coded partner ($27.4\%$) than in the untrustworthy case ($12.9\%$).
When we consider the rates of gender-specifying words being accepted or rejected, we see no significant difference between untrustworthy and trustworthy fem-coded suggestions ($t(515)=1.364$, $p_{\text{FDR}}\approx 0.3400$, $d=0.12$). Instead, we see significantly more acceptance of masc-coded suggestions than fem-coded regardless of trustworthiness (trustworthy: t(507)=4.224t(507)=-4.224, $p_{\text{FDR}}\approx 0.0003$, d=0.375d=-0.375; untrustworthy: t(325)=7.066t(325)=-7.066, $p_{\text{FDR}}<0.0001$, d=0.913d=-0.913) and significantly more acceptance of untrustworthy masc-coded suggestions than trustworthy ($t(317)=4.384$, $p_{\text{FDR}}\approx 0.0002$, $d=0.569$ ).
Overall, we see that the rejection rate results concur with our and cao-etal-2022-theory’s findings about people’s gender-trustworthiness stereotypes and support H2b. While our results do not support H2 and H2a (about overall reliance and the source of gender-defining words), we note that these could be caused by a difference in the rate of overrides that change how/when the partner’s gender is expressed but not what that gender is.
@@ 6.3. Scenario: Wedding
In this scenario, we consider a wedding between two partners. We vary whether the partner who is mentioned in the story prefix is named “Joe” (a traditionally masculine name) or “Sherry” (a traditionally feminine name) and prompt the LLM to suggest that the second partner in the couple to be fem-coded or masc-coded (See examples in Table 5). Participants specified the second partner character’s gender in about $89\%$ of stories (in both treatment and control conditions). We treat the pro-stereotypical conditions to be those where the genders of the partners are suggested to be different and the anti-stereotypical to be those where they are suggested to be the same.
@@
Bar charts showing the gender distribution of second partners under varied suggestions, grouped by whether the first partner is named “Joe” or “Sherry.” When the first partner is “Joe,” the charts show more feminine-coded second partners and fewer queer relationships, even with queer suggestions. When the first partner is “Sherry,” the charts show higher acceptance of queer relationship suggestions.
@@
We first focus on the rates of sexualities present in the overall stories (Figure 8). We start by considering what kinds of suggestions yield similar or different stories to those written without suggestions. When the first partner is “Sherry” and the model suggests a queer relationship (i.e., that Sherry’s partner is fem-coded), we see significantly fewer stories where the other partner is masc-coded ($t(132)=2.973$, $p_{\text{FDR}}\approx 0.0172$, $d=0.547$) and significantly more stories where the other partner is fem-coded (t(132)=2.695t(132)=-2.695, $p_{\text{FDR}}\approx 0.0328$, d=0.496d=-0.496). There are no significant changes to the distribution of Sherry’s partners’ gender when straight suggestions are provided (fem-coded partner: $t(124)=0.657$, $p_{\text{FDR}}\approx 0.7019$, $d=0.123$; masc-coded partner t(124)=0.422t(124)=-0.422, $p_{\text{FDR}}\approx 0.8234$, d=0.079d=-0.079). The direction of these trends are mirrored for “Joe” stories but with insignificant changes. That is, “Joe” is written with masc-coded partners more often (t(115)=1.688t(115)=-1.688, $p_{\text{FDR}}\approx 0.2218$, d=0.361d=-0.361) and fem-coded partners less often when given queer suggestions ($t(115)=2.143$, $p_{\text{FDR}}\approx 0.1071$, $d=0.459$), and there are no apparent changes when given straight suggestions to the rate of fem-coded (t(104)=0.556t(104)=-0.556, $p_{\text{FDR}}\approx 0.7514$, d=0.121d=-0.121) or masc-coded partners (t(104)=0.154t(104)=-0.154, $p_{\text{FDR}}\approx 0.9479$, d=0.034d=-0.034). These findings suggest that indeed participants’ default behavior without suggestions is more similar to the “straight” conditions than the “queer” ones, though participants seem to be more resistant to accepting “queer” suggestions for “Joe” than “Sherry”.
@@
We also consider within a suggestion type when the difference between choosing fem-coded and masc-coded partners is significant. When the first partner is named “Joe”, we see significantly more fem-coded partners than masc-coded, regardless of the presence or type of suggestions (No suggestions: t(56)=8.825t(56)=-8.825, $p_{\text{FDR}}<0.0001$, d=2.318d=-2.318; queer suggestions: t(174)=6.035t(174)=-6.035, $p_{\text{FDR}}<0.0001$, d=0.91d=-0.91; straight suggestions: t(152)=16.063t(152)=-16.063, $p_{\text{FDR}}<0.0001$, d=2.589d=-2.589). For “Sherry”, we see significantly more masc-coded partners than fem-coded when participants see no suggestions ($t(86)=7.704$, $p_{\text{FDR}}<0.0001$, $d=1.643$) or straight suggestions ($t(162)=12.859$, $p_{\text{FDR}}<0.0001$, $d=2.008$). We see a trend in the same direction with queer suggestions, but here it is not significant ($t(178)=2.12$, $p_{\text{FDR}}\approx 0.1097$, $d=0.316$). These findings again imply that without suggestions, participants default to hetero-normative stories and continue to write them when prompted (H1). When queer stories are suggested, depending on the gender of the partner that is fixed in the story prefix, participants may start to accept more queer stories, but continue to prefer to write hetero-normative stories overall.
@@
Based on these results, even when writing with a “perfectly debiased” predictive text system that, for instance, has no preference for gender pairings, we would expect to continue to see far more straight stories than queer ones.
@@
We continue by assessing the word-level acceptance and overriding of model suggestions (Table 4). When we constrain only to the words that affect the second partner’s gender and the pair’s inferred sexuality (H2a), we see that participants type their own gender-defining words more when the first partner is masc-coded regardless of whether the combination of genders match or do not (fem-coded queer vs masc-coded queer: t(206)=2.566t(206)=-2.566, $p_{\text{FDR}}\approx 0.0424$, d=0.358d=-0.358; fem-coded straight vs masc-coded straight: t(195)=3.554t(195)=-3.554, $p_{\text{FDR}}\approx 0.0031$, d=0.507d=-0.507). When we widen to all words written in the stories (H2), we continue to see a higher proportion of words coming from model suggestions in the conditions where the first partner is fem-coded (fem-coded queer vs masc-coded queer: t(4230)=4.019t(4230)=-4.019, $p_{\text{FDR}}\approx 0.0005$, d=0.124d=-0.124; fem-coded straight vs masc-coded straight: t(3699)=4.373t(3699)=-4.373, $p_{\text{FDR}}\approx 0.0001$, d=0.144d=-0.144), but now we also see we see marginally more acceptance of model suggestions in the condition where the fem-coded partner is suggested to be marrying a masc-coded character ($t(4010)=2.381$, $p_{\text{FDR}}\approx 0.0610$, $d=0.075$). Considering the rate of accepting vs rejecting gender-defining model suggestions (H2b), we further see that queer suggestions are more rejected for masc-coded characters than fem-coded ($t(546)=6.786$, $p_{\text{FDR}}<0.0001$, $d=0.597$) and that for masc-coded characters, queer suggestions are rejected significantly more than straight suggestions (t(607)=2.901t(607)=-2.901, $p_{\text{FDR}}\approx 0.0187$, d=0.237d=-0.237).
6.4. Scenario: President
In this scenario, the story prefix describes a new president laying out their policy vision (“In the first 100 days, the new president was determined to focus”). The model is prompted (unseen by the participant) to suggest that the president is of a particular gender (man or woman) and communion (“benevolent” or “threatening”). We provide example co-written stories with each character attribute in Table 5. The in-context examples given to the model focused on the benevolence axis as it applies to foreign policy, but the final written stories discuss a wider array of policy areas. In both cao-etal-2022-theory and our post-survey, participants viewed men as less benevolent or warm than women (See Table 16).
6.4.1. Effects on Gender Alone
First, we consider how suggestions affect how participants specify a president character’s gender. Participants specified the president character’s gender in about $57\%$ of stories (and $55\%$ of co-written stories).
At the Story Level
@@ 6.4.1. Effects on Gender Alone
Bar charts showing the gender distribution of presidents under no-suggestion, masculine-coded, and feminine-coded suggestion conditions, further grouped by benevolence. Masculine-coded suggestions produce a similar distribution to no suggestions, while feminine-coded suggestions increase feminine-coded presidents without exceeding masculine-coded. Across conditions, most presidents are labeled benevolent regardless of suggestions.
@@
We first look into how suggestions affect the president character’s gender in the overall stories (9(a)). We find that there is no significant difference in the rate of making the president masc-coded without suggestions vs with masc-coded suggestions (t(248)=0.137t(248)=-0.137, $p_{\text{FDR}}\approx 0.9488$, d=0.019d=-0.019) and similarly for making the president fem-coded with masc-coded suggestions (t(248)=1.106t(248)=-1.106, $p_{\text{FDR}}\approx 0.4750$, d=0.153d=-0.153). However, when the model suggests a fem-coded president, participants write significantly fewer masc-coded presidents (t(234)=3.576t(234)=-3.576, $p_{\text{FDR}}\approx 0.0029$, d=0.502d=-0.502) and marginally more fem-coded presidents ($t(234)=2.429$, $p_{\text{FDR}}\approx 0.0565$, $d=0.341$) than those who did not receive suggestions. Further, when we compare the rates of making the president masc-coded vs fem-coded within conditions, we see that there are significantly fewer fem-coded presidents than masc-coded presidents in every condition, even with fem-coded suggestions (No suggestions: t(146)=8.947t(146)=-8.947, $p_{\text{FDR}}<0.0001$, d=1.471d=-1.471; masc-coded suggestions: t(350)=14.755t(350)=-14.755, $p_{\text{FDR}}<0.0001$, d=1.573d=-1.573; fem-coded suggestions: t(322)=4.343t(322)=-4.343, $p_{\text{FDR}}\approx 0.0002$, d=0.483d=-0.483).
@@
These results support H1, namely that the distribution of the gender of the president characters when people write by default without suggestions is more similar to the distribution when people write with masc-coded suggestions than fem-coded suggestions. Based on these findings, we would expect stories written with a “debiased” predictive text model (that suggests fem-coded presidents less often than in our experiment) to still yield significantly more stories with masc-coded presidents than fem-coded presidents.
@@
Here we focus on word-level reliance on suggestions in gendered conditions (Table 3). We find participants’ overall reliance on model suggestions is not affected by the which gender is suggested (H2; t(6734)=0.979t(6734)=-0.979, $p_{\text{FDR}}\approx 0.5407$, d=0.024d=-0.024). However, when we only consider the story words that specify the president character’s gender, we see a lower rate of overrides or edits in the masc-coded president settings (H2a; $t(203)=5.269$, $p_{\text{FDR}}<0.0001$, $d=0.742$). Similarly, when we consider only model suggestions that would specify the president character’s gender, we see a significantly higher rejection rate for words that describe the president as fem-coded (H2b; $t(836)=6.362$, $p_{\text{FDR}}<0.0001$, $d=0.442$). These results support H2a-b as participants are more likely to accept model suggestions of masc-coded presidents than fem-coded presidents.